### v1

# Large Language Models for Orchestrating Bimanual Robots

Abstract
Although there has been rapid progress in endowing robots with the ability to solve complex manipulation tasks, generating control policies for bimanual robots to solve tasks involving two hands is still challenging because of the difficulties in effective temporal and spatial coordination. With emergent abilities in terms of step-by-step reasoning and in-context learning, Large Language Models (LLMs) have taken control of a variety of robotic tasks. However, the nature of language communication via a single sequence of discrete symbols makes LLM-based coordination in continuous space a particular challenge for bimanual tasks. To tackle this challenge for the first time by an LLM, we present LAnguage-model-based Bimanual ORchestration (LABOR), an agent utilizing an LLM to analyze task configurations and devise coordination control policies for addressing long-horizon bimanual tasks. In the simulated environment, the LABOR agent is evaluated through several everyday tasks on the NICOL humanoid robot. Reported success rates indicate that overall coordination efficiency is close to optimal performance, while the analysis of failure causes, classified into spatial and temporal coordination and skill selection, shows that these vary over tasks. The project website can be found at http://labor-agent.github.io
I Introduction
While humans routinely operate with both hands in their daily lives, achieving effective bimanual coordination remains a formidable challenge in robotics as the dual-arm setting faces complexity in terms of temporal and spatial coordination. Bimanual manipulation offers a significant difficulty for the existing planning-based and learning-based methods. Previous planning-based methods [1, 2] largely focus on motion planning for coordinating two arms through hand-designed primitives or predefined trajectories, limiting their application in dynamic or complex task spaces. On the other hand, learning-based methods, such as Reinforcement Learning (RL) and Imitation Learning (IL), enable a robot to learn control policies from human-designed rewards [3, 4] or human-teleoperated demonstrations [5, 6]. However, it is notoriously known that RL algorithms are hard to train, especially in dual-arm settings that have high degrees of freedom (DoFs), and collecting human demonstrations for bimanual robots is non-trivial. At the same time, the high complexity associated with the variety of bimanual patterns suggests that low-level control policies and high-level planning must be considered for an integrated control system design [7].
Large Language Models (LLMs) have demonstrated remarkable knowledge and reasoning capabilities [8, 9], leading to a revolutionary change in numerous fields [10]. In robotics, an LLM is typically employed as a high-level planner to reason and invoke primitive skills for accomplishing complex tasks described in natural language [11, 12, 13]. However, as far as we know, all previous methods focus on unimanual manipulations with complex chains of skills and lack explicit analysis of bimanual tasks. In this research, we attempt to further explore the feasibility of leveraging LLMs in robotic bimanual control, especially focusing on high-level temporal and spatial coordination between two hands. Our research question is: How can LLMs be employed to achieve versatile bimanual manipulation control in diverse robotic tasks?
Figure 1: Illustration of the LABOR agent. During the execution of the task, with the designed prompt and skill library, the LLM generates the plan step by step, including a sequential or simultaneous control command for the left and right hand. The bimanual robot then performs actions on the environment according to the command and the results get feedback to the LLM for the next action, and so on until the end of the task.
To this aim, we introduce the LABOR (LAnguage-model-based Bimanual ORchestration) agent, which uses an LLM to effectively coordinate bimanual control to deal with complex manipulation tasks, as shown in Fig. 1. Two types of control patterns, i.e., sequential and simultaneous control, are explicitly predicted by the LLM, given a designed prompt with a set of possible skills. With a high-level control timeline, a bimanual robot is able to complete a long-horizon complex task by combining different control policies through the cooperation of both hands at different stages. We evaluate the LABOR agent on the semi-humanoid bimanual NICOL, the Neuro-Inspired COLlaborator [14], on three typical everyday tasks. The experimental results show that the LABOR agent is effective in coordinating both hands in long-horizon tasks.
II Related Work
In this section, we describe the related work on planning-based and learning-based methods for bimanual manipulations, and LLM’s use in robotics.
II-A Bimanual Manipulation
II-A1 Planning-based methods
Early work [15, 7] mainly focuses on planning algorithms for dual-arm settings under the closed-chain constraint, which rely on static known dynamic models to generate kinematic configurations that are suitable for two arms. With engineered primitives or predefined trajectories [1, 16, 2, 17], bimanual robots exhibit some flexibility when solving tasks like folding shirts or scooping food. However, such planning-based methods are not adaptive to dynamic environments and mainly rely on task-specific primitives, which are costly to design or collect in a more complex task setting.
II-A2 Learning-based methods
Some works have employed RL [18, 3, 4, 19] and IL [5, 20, 21, 6] framework in dual-arm settings. These typically require massive samples and training time, due to the high dimensionality of the action space in a bimanual setting. Although some recent works proposed to design parameterized primitives to reduce the size of the action space for exploration, these movement primitives normally require costly hard-coded motion design or human-teleoperated demonstration collection, limiting their applications in complex environments.
Instead of designing many task-specific primitives, we use easily designed general primitives for the robot and attempt to utilize LLMs to generate chains of skills to accomplish complex bimanual motions. Furthermore, the approaches listed above do little to analyze the properties of the task at a higher level, but rather adopt a single control policy for controlling the robot to accomplish the task. In this sense, the control policy in their tasks is normally fixed and monolithic which tackles only simple and short-duration tasks. In contrast, we explore the usage of LLMs to generate control policies from a higher perspective that considers the current stage of the task at hand, thus adapting to complex and variable long-horizon tasks.
II-B LLM in Robotics
Learning robotic control policies to address various manipulation tasks often requires the collection of massive datasets [22, 23, 24, 25]. Recent works have leveraged common sense and established knowledge from LLMs to control robots in tackling complex tasks. Some studies have delved into distilling LLM expertise into localized models. For instance, researchers have utilized LLMs’ programming capabilities to generate executable code for fulfilling open instructions [26, 27, 28], or to provide direct guidance for RL policies [29, 30]. Other approaches involve decomposing complex tasks into manageable sub-tasks using LLMs, thus enabling robots to address them individually with pre-defined skills [12, 11], i.e., language-conditioned policies. We adopt the latter paradigm to construct a versatile robot equipped with atomic skills that can be efficiently reused across multiple tasks or potentially enhanced for new abilities [13]. Specifically, our robot is furnished with general skills for both arms, offering a vast number of potential combinations. However, this inherent complexity of combinatorial capabilities poses challenges to efficient reasoning and planning in bimanual manipulation, which will be explored in this study.
III Method
This section introduces the LABOR agent regarding two main aspects, Bimanual Manipulation, which describes the control policies adopted, and LLM Orchestration, which describes how the control policies can be coordinated using an LLM.
III-A Bimanual Manipulation
As described in earlier works [5, 31, 32], basic bimanual manipulation control mainly involves the consideration of coordination and, when required, synchronization. For example, grasping or releasing small objects does not need bimanual coordination as one single hand can accomplish such a task. In contrast, to lift a large object like a bowl, two hands are needed to hold it, and they must coordinate symmetrically. In the task of opening a pair of scissors, two hands need to cooperate asymmetrically, for example, one hand needs to stabilize the scissors first, and another hand slides the handle away from the other handle.
Figure 2: Control policies adopted by the LABOR agent.
Considering the spatial and temporal constraints, we adopt two types of control policies for two hands, simultaneous control for symmetric processes, and sequential control for uncoordinated and asymmetric processes, as shown in Fig. 2. In a macroscopic view of long-horizon bimanual manipulation tasks, a combination of processes can be involved. For example, imagine that there is an apple and a bowl on the table, the task is placing the apple in the bowl and holding the bowl to a serving position. There will be uncoordinated processes such as grasping and releasing the apple, and then symmetrically coordinated processes such as joint holding and moving are required. During the task accomplishment, which control policy to adopt should be decided first, based on the current spatial and temporal situations, and then the specific control commands to coordinate two hands.
III-B LLM Orchestration
Figure 3: Prompts for the LLM to orchestrate a bimanual robot to accomplish the task. The general prompting template, indicated in gray, remains consistent across various tasks, while task-specific details such as task description and manipulation rules are presented in standard black text.
An LLM exhibits remarkable abilities in terms of step-by-step reasoning, in-context- and instruction-following responses in robotics, which makes it possible to generate policies under temporal and spatial constraints in bimanual manipulations. As shown in Fig 1, the presented LABOR agent involves human user inputs like the initial prompt and skill library, the LLM coordinator, and the bimanual robot executor. As depicted in Fig. 3, the designed prompt contains background information regarding the bimanual robot, like coordinates and workspaces of two hands, following some rules specifying the bimanual control principles when facing different situations. At the same time, the skill library is available to the LLM in the form of tool APIs. Given this, the prompts explicitly require the LLM to generate policies that are sequential or simultaneous, with specific control commands based on the provided skill library. The skill library is designed based on control policies and consists of two main types: single-handed skills and both-handed skills. After executing an action assigned by the command, the LLM obtains observations from the execution results, e.g., an action was completed by the left or right hand, the skill execution failed because the assigned action was not in the valid workspace of the selected hand, etc. Iterating in this way, the LLM progressively generates a solution to the task while self-correcting mistaken commands, thus generating a chain of skills throughout the lifespan of the task, as shown in Fig. 4.
Figure 4: An illustrative example of the chain of skills with sequential (seq.) or simultaneous (sim.) control policy. Commands generated for each timestep follow one specific control policy, and the skills contained are chosen from the designed skill library.
In summary, the LABOR agent uses the LLM to explicitly coordinate the operation of both hands while continuously generating skill-based commands, to accomplish complex long-horizon bimanual tasks effectively.
IV Experiments
To determine the extent to which an LLM can manage to control the bimanual robot in everyday tasks, we deploy the LABOR agent on the semi-humanoid NICOL [^1] The NICOL robot website can be found at https://www.inf.uni-hamburg.de/en/inst/ab/wtm/research/neurobotics/nicol.html, the Neuro-Inspired COLlaborator [14], on three typical everyday tasks in the CoppeliaSim [^2] https://www.coppeliarobotics.com/ simulation. The implementation of the LABOR agent and experiment design details will be open-sourced upon the paper’s publication.
IV-A LABOR Setup
For the NICOL robot, we design a skill library [^3] The skills are developed based on the nicol_coppeliasim Python library, https://github.com/knowledgetechnologyuhh/nicol_coppeliasim as listed in Table I. There are three categories of skills: Single, Both, and Info. Single or Both denotes that the skill is designed for a single hand or both hands, while Info refers to those functions that can obtain information about the robot and the environment.
We use the OpenAI GPT models [^4] We chose two models, gpt-4-0125-preview and gpt-3.5-turbo-0125, for our experiments. as the LLM for inference in the LABOR agent. Since the accomplishment of bimanual tasks requires the LLM to generate sequential or simultaneous control several times, we take the LangChain [^5] https://github.com/langchain-ai/langchain Python library framework to enable the LLM to generate a chain of skills step by step. Based on the LangChain tool design structure, several tools are provided to the LLM as follows:
  • sequential control: llm_seq_control
    required inputs: (left_hand_command, left_parameter, right_hand_command, right_parameter, if_left_first)
  • simultaneous control: llm_simul_control
    required inputs: (both_hand_command, parameter)
  • get information: (get_info_command, parameter)
where left_hand_command and right_hand_command are chosen from the “Single" skill category with the parameters specified in left_parameter and right_parameter respectively, if_left_first is a bool value that specifies if the left_hand_command should be executed at first or not, and both_hand_command and get_info_command are chosen from Both and Info skill categories respectively. With tools for different control policies, the LABOR agent enables the LLM to choose sequential or simultaneous control in a high-level manner explicitly.
TABLE I: The full list of skills used for Nicol robot
Skill Name Parameters Catagory
move_to_pose (side, obj_name, offset) Single
move_to_jab_at (side, obj_name, direction)
move_to_press_on (side, obj_name)
delta_move (side, direction, distance)
top_grasp (side, obj_name)
side_grasp (side, obj_name)
release (side)
flip_down (side)
wait (side)
move_both_to_pose (obj_name) Both
lateral_hold (obj_name)
reset_both None
get_arm_state None Info
get_object_distance (obj_A, obj_B)
get_joint_angle (joint)
IV-B Robot Setup
NICOL has a head that can express stylized facial expressions, and two Robotic Open-Manipulator-Pro arms with adult-sized five-fingered Seed Robotics RH8D hands (for details see [14]). With such equipment, NICOL has a large bimanual workspace to handle everyday objects and tasks. NICOL’s workspace with objects in the real and simulated world is shown in Fig 5.
Based on the objects listed, there are three tasks designed in the experiments: ControlScissors, ServeWater, and HoldBowl, which all require the cooperation of two hands with a chain of skills to accomplish.
(a) real world
(b) simulated world
Figure 5: NICOL workspace with everyday objects. In the simulated environment, objects placed from left to right are apple, banana, blue cup, bowl, scissors, and yellow cup.
(a) ControlScissors
(b) ServeWater
(c) HoldBowl
Figure 6: Description of bimanual tasks designed for the NICOL robot. There are critical stages on three tasks illustrated: (a) the left hand presses on the scissors handle to stabilize the scissor while the right hand slides another handle using its finger to open the scissor; (b) the right hand flips down the water-filled cup to pour out the ‘water’ (small ball) to the empty cup while the left hand is holding the grasped empty cup; (c) the left and right hand is holding together the bowl with the apple and banana inside.
(a) ControlScissors
(b) ServeWater
(c) HoldBowl
Figure 7: Illustration of the number of timesteps used when accomplishing three tasks in successful cases. The vertical coordinate represents the number of timesteps required for all successful cases, and the corresponding horizontal length of each bar denotes the count of cases that required that number of timesteps. In ServeWater and HoldBowl tasks, orange color denotes scenes where objects are initially posed at different sides (left/right) of the robot, while blue denotes all objects are placed on the same side. The dashed lines of the corresponding colors indicate the minimum number of timesteps required, as given by the human-designed solution.
In the ControlScissors task, there is a pair of scissors placed horizontally on the work table, with its tip pointing in various directions: Front, Behind, Left, and Right. The goal of the task is to stabilize the handle on one side while sliding the handle on the other side in a certain direction to Open or Close the scissors. In this way, eight variants have been produced, depending on the combination of the direction in which the tips of the scissors are pointed and the task requirements of the switch. The difficulty of this task comes from the fact that the LLM needs to understand the spatial relationship between the two handles, and then reason the direction of the handle’s movement after stabilizing the scissors.
In the ServeWater task, there are two cups on the work table. The yellow cup is empty and the blue cup is water-filled (we use a small ball as simulated water). The goal of the task is to use the yellow cup to serve water at a specified serving position. This means, that the water has to be transferred from the blue cup to the yellow cup, the yellow cup is then brought to the serving position, and the blue cup is released back to its original position. The cups can be in various spatial relations including, including Both_Right, Both_Left, Left_Blue_Right_Yellow, and Left_Yellow_Right_Blue. Since the robot cannot reach with its right arm over the left side of the table, and vice versa, different spatial configurations afford different task solutions. However, this relationship information is not passed explicitly to the LLM, which instead receives raw information about the coordinates of the two cups, placed on the table at random positions. The task itself requires a long skill chain to accomplish and coupled with the changing positional relationship between cups and hands, choosing appropriate control policies effectively is even more challenging.
In the HoldBowl task, there is one large bowl, one apple, and one banana on the work table. The goal of the task is to grasp the apple and the banana, put them into the bowl, and hold the bowl to a specified serving position. Similar to ServeWater, the positions of the apple and banana bring complexity and variety to this task, including Both_Right, Both_Left, Left_Apple_Right_Banana, and Left_Banana_Right_Apple, as well as the choices of control policies. For each case, the sequential skills should be adopted first to transfer the objects into the bowl, and then the simultaneous ones for holding the bowl to move to the specified position.
IV-C Results
In this section, we report the results of the LABOR agent in controlling the NICOL robot to solve the above tasks, including success rate and failure analysis. To evaluate the effectiveness, we repeat each task 100 times and for each time a variant of the task is randomly chosen. The success rates and failure rates in three aspects of LABOR when using GPT-4 as the LLM on these tasks are listed in Table II.
TABLE II: Success and Failure rates of NICOL on three everyday tasks
Task ControlScissors ServeWater HoldBowl
Success 79% 89% 86%
Failure Spatial 12% 5% 4%
Temporal 1% 6% 1%
Skill 8% 0% 9%
It is observed that the LABOR agent with GPT-4 demonstrates extraordinary performance in orchestrating bimanual control in long-horizon manipulation tasks. Regarding the properties of bimanual manipulation and the LLM itself, we found that the failure cases mainly come from three aspects including Temporal Coordination, Spatial Coordination, and Skill Association. In this sense, Temporal failure refers to the fact that the LLM fails to generate skills in the correct order, Spatial failure refers that the LLM mishandles the spatial relationships in the operation, such as the relationship between two objects or between the robot’s hand and an object, and Skill failure refers to the understanding deviance of the LLM in the correlation between the designed skills and the task requirements. To simplify the analysis process, failures are counted based on the error type that occurred the first time.
It was observed that GPT-4 encountered some difficulties with spatial reasoning (i.e., moving in a certain direction to open or close the scissors) in the ControlScissors task. Whereas, the main errors in the ServeWater task came from temporal and spatial coordination, such as executing pouring before moving the blue cup over the yellow cup, executing grasping when the selected hand had already grasped another object, executing grasping when the object that was outside of the selected hand’s workspace, etc. Moreover, some skill comprehension biases happened in the ControlScissors and HoldBowl tasks, such as moving the scissors handle with the hand currently pressing on it, executing grasping the bowl with a single hand, and so on.
In contrast, when implementing LABOR with GPT-3.5, it only achieved a 49% success rate in the ControlScisssor task, which might be due to the task requiring fewer planning steps. Apart from that, it demonstrates its inability to accomplish either ServeWater or HoldBowl tasks. The failures observed in these tasks couldn’t be neatly categorized into three aspects, as in such long-horizon tasks, the GPT-3.5 model’s understanding of the skills and its temporal and spatial coordination was incorrect at the first step. In addition, it is not sensitive enough to the information feedback by the environment, causing it to stay stuck in repeating wrong commands.
In addition to the success rate, we also counted the lengths of the skill chains generated by GPT-4 in all cases where the task was successfully solved, and compared them with the optimal human-designed skill chain lengths, as shown in Fig. 7. Since there are variants in the ServeWater and HoldBowl quests that single-hand sequential operation, the statistics in these two tasks are reported separately based on the position in which the two objects are located, i.e., Same Side and Two Sides. The number of timesteps indicated by the dashed lines is considered to be the optimal lower bound for the number of timesteps required. We can observe that there are many cases in ServeWater and ControlScissors tasks, where an optimal skill chain can be achieved that is comparable to the human level, while it is pretty close in the HoldBowl task. To ensure the safety of the two-handed operation, the GPT-4 generates some waits for single hands thus bringing about an increase in the length of the chain. At the same time, humans understand the task well enough, thus allowing for the most efficient utilization of both hands. Notably, due to GPT-4’s strong in-context reasoning and understanding ability, sometimes long skill chains can still ensure task accomplishment in the ServeWater and HoldBowl tasks.
In summary, the LABOR agent with GPT-4 shows remarkable performance in coordinating the NICOL robot in long-horizon tasks, not only in terms of the success rate but also in outputting skill chains comparable to the human level.
V Conclusions
In this work, we introduce LABOR, an agent that utilizes an LLM to effectively orchestrate bimanual control when performing robotic manipulation tasks. Experimental results on the NICOL robot for three typical everyday tasks showcase the effectiveness of the LABOR agent.
Limitations and Future Work. Similar to other LLM-based methods, the LABOR agent is dependent on the performance of LLM like the GPT-4 model, requiring its reasoning and understanding of the correlations of primitive skills and task solutions. Besides, since the agent is currently solely running in the simulated environment instead of the real world, some potential differences in environmental dynamics might not be sufficiently analyzed or considered, which brings some unknowns in transferring the agent. In future work, the learning-based skills combined with vision-based models will be developed to reinforce the abilities of LABOR agent in bimanual manipulation. As skills expand, the LABOR agent should be able to control bimanual robots to perform more complex tasks. Furthermore, it is a promising direction to include obstacle detection and collision-free motion planning models in dealing with rich-contact tasks.
VI ACKNOWLEDGMENT
The authors would like to thank Jan-Gerrit Habekost and Lennart Clasmeier for their work and support on the NICOL robot in the simulation.